# Spectral Convolutional Conditional Neural Processes

### v5

@@ 1 Introduction
Stochastic processes offer a mathematical framework for modeling systems that evolve with inherent randomness over continuous domains such as time and space. They underpin a wide range of scientific applications—from spatio-temporal climate dynamics to biological and physical systems—thereby motivating the development of machine learning methods that can learn from data generated by such phenomena (mathieu2021contrastive; vaughan2021convolutional; allen2025end; ashman2025gridded; dupont2021generative). Among classical approaches, Gaussian processes (GPs; rasmussen2006gaussian) provide a Bayesian framework with closed-form inference and uncertainty quantification. However, their cubic computational cost from matrix inversion and the difficulty of specifying suitable kernels—especially in high-dimensional settings—limit their scalability.
@@
Motivated by the success of deep neural networks in large-scale function approximation, neural network-based alternatives have emerged. Neural processes (NPs; garnelo2018conditional; garnelo2018neural) exemplify this paradigm, combining ideas from GPs and deep learning within a meta-learning framework. By exposing the model to multiple realizations of an underlying stochastic process, each treated as a distinct task, NPs learn shared structures across tasks to parameterize a neural mapping that directly gives an approximation of the corresponding posterior predictive distribution (bruinsma2024convolutional). Once trained, the model enables efficient probabilistic predictions on new tasks without further training.
@@
Since the introduction of conditional neural processes (CNPs, garnelo2018conditional) as the first class within the NPs family (jha2022neural), numerous extensions have been proposed to enhance their effectiveness. One prominent line of work focuses on incorporating explicit inductive biases into CNPs in order to better capture the symmetries that commonly arise in scientific applications (gordon2019convolutional; kawano2021group; holderrieth2021equivariant; huang2023practical; ashman2024translation; ashman2024approximately). Another major direction seeks to move beyond the mean-field factorized Gaussian predictive distributions to which CNPs are limited. A widely adopted approach augments CNPs with stochastic latent variables, giving rise to the family of latent neural processes (LNPs; garnelo2018neural; louizos2019functional; wang2020doubly; foong2020meta; lee2020bootstrapping; volpp2021bayesian; wang2022bridge; wang2022learning; kim2022neural; jung2023bayesian; lee2023martingale; xu2023deep). Complementary efforts explore explore autoregressive prediction schemes (bruinsma2023autoregressive; nguyen2022transformer), Gaussian predictive distributions with non-diagonal covariances (bruinsma2021gaussian; markou2022practical), and quantile-based parameterizations of the predictive distribution (mohseni2023adaptive).
@@
This work focuses on CNPs, particularly convolutional CNPs (ConvCNPs; gordon2019convolutional), which were the first to endow NPs with translation equivariance. ConvCNPs introduce the convolutional deep set construction which characterizes a broad class of translation-equivariant mappings over finite, potentially unstructured sets of observations as a composition of a functional embedding with a translation-equivariant operator, typically realized through convolutional neural networks (CNNs; fukushima1980neocognitron; lecun1989backpropagation; lecun1998gradient).
@@
Despite their effectiveness, ConvCNPs can struggle to aggregate information from observations spread across large spatial domains—a challenge that becomes particularly pronounced in sparse data regimes. This limitation stems from their reliance on local convolutional kernels with small receptive fields, which hampers their ability to model long-range dependencies. A natural remedy is to enlarge the kernel size to extend the receptive field; however, this approach rapidly increases the number of model parameters and computational cost (romero2021ckconv). Alternatively, transformer-based architectures can capture long-range interactions but incur quadratic, rather than linear, computational complexity in the number of observations (vaswani2017attention; nguyen2022transformer).
@@
In this work, we pursue an alternative paradigm that represents functions in the frequency domain, inspired by the well-established observation that many natural processes exhibit energy concentration in low-frequency bands (field1987relations; ruderman1993statistics; wainwright1999scale). This property allows for efficient approximation using only a subset of dominant spectral coefficients, enabling tractable computations while preserving the signal’s global structure. By parameterizing convolution kernels directly in the Fourier domain over a finite set of frequencies and leveraging the convolution theorem, we can attain large effective receptive fields without incurring prohibitive computational costs.
@@
While spectral methods have been extensively studied in neural operator learning for partial differential equations (PDEs, chen1995universal; kovachki2021universal; kovachki2023neural), their application within NPs framework remains relatively unexplored. To bridge this gap, we propose spectral convolutional conditional neural processes (SConvCNPs)—a class of models that adopt Fourier neural operators (FNOs; li2020fourier) to realize global spectral convolution while maintaining computational efficiency. Across a suite of synthetic and real-world benchmarks, SConvCNPs perform competitively with state-of-the-art baselines, illustrating how ideas from neural operators can enhance the flexibility and performance of NPs.
@@ 2.1 Fourier Neural Operators
Neural operators (chen1995universal; li2020fourier; kovachki2023neural) are neural network architectures designed to learn mappings between function spaces rather than finite-dimensional vectors. Much like conventional feed-forward networks, they consist of stacked layers that alternate between operator-based transformations and pointwise nonlinearities. The common transformation choice is an integral operator with kernel $\kappa:\mathcal{X}\times\mathcal{X}\to\mathcal{Y}$, acting on an input function $v:\mathcal{X}\to\mathcal{Y}$ as
$\mathcal{K}[v](x)=\int\kappa(x,s)\,v(s)\,ds,$
@@
where $\mathcal{X}={\mathbb{R}}^{d_{x}}$ and $\mathcal{Y}={\mathbb{R}}^{d_{y}}$ are Euclidean spaces with dimensions $d_{x}$} and $d_{y}$}, respectively. While this work focuses on operators that are linear in $v$, nonlinear formulations have also been explored, including continuous formulations of softmax attention (ashman2024translation; calvello2024continuum). When the kernel is stationary, i.e. $\kappa\left(x,s\right)=\kappa\left(x-s\right)$, the operator simplifies to a convolution, $\mathcal{K}[v]=\kappa\ast v$, connecting neural operators to CNNs, where $\kappa$ is parameterized by learnable weights.
@@
Typically, convolutional kernels are spatially local and possess a limited receptive field (luo2016understanding; Peng_2017_CVPR; wang2018non). Consequently, modeling long-range dependencies requires increasingly large kernels, which substantially inflates the parameter count. The Fourier Neural Operator (FNO, li2020fourier) overcomes this limitation by exploiting the convolution theorem (bracewell1966fourier; oppenheim1999discrete), which re-expresses convolution as:
$\mathcal{K}[v](x)=\mathcal{F}^{-1}\Big[\mathcal{F}[\kappa](\xi)\cdot\mathcal{F}[v](\xi)\Big](x),$ (1)
@@
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively. Rather than parameterizing the kernel $\kappa$ in the physical domain, the FNO learns its representation directly in the Fourier domain. When the input function $v$ is approximately band-limited—i.e., its spectrum $\mathcal{F}[v](\xi)$ carries negligible energy for frequencies $\lVert\xi\rVert>\xi_{0}$—the high-frequency components can be truncated with minimal loss of information. This property, observed in many natural signals (field1987relations; ruderman1993statistics; wainwright1999scale), allows setting $\mathcal{F}[\kappa](\xi)=0$ outside the retained frequency band without significant information loss.
@@
2.2 Neural Processes
Let $\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$ denote the set of probability measures over functions $f:\mathcal{X}\to\mathcal{Y}$. We assume a data-generating process $\mu\in\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$, from which functions are sampled. A task $\mathcal{D}$ is defined as a finite collection of input–output pairs drawn from a noisy realization $f\sim\mu$, partitioned into a context set and a query [^2] The NPs literature conventionally uses the term “target points.” To avoid notational ambiguity with time indices, we instead use “query points.” set (bruinsma2024convolutional; ashman2024translation; ashman2024approximately):
$\mathcal{D}=\left(\mathcal{D}_{c},\mathcal{D}_{q}\right)=\left(\left\{\left(x_{c,k},y_{c,k}\right)\right\}_{k\in\mathcal{I}_{c}},\;\left\{\left(x_{q,k},y_{q,k}\right)\right\}_{k\in\mathcal{I}_{q}}\right),\vskip-7.11317pt$ (2)
where
$y_{c,k}=f(x_{c,k})+\epsilon_{c,k},\quad y_{q,k}=f(x_{q,k})+\epsilon_{q,k},\quad\epsilon_{c,k},\epsilon_{q,k}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma_{0}^{2}).$
Here, $\sigma_{0}>0$ denotes the observation noise scale, and $\mathcal{I}_{c},\mathcal{I}_{q}\subset\mathbb{N}$ index the context and query sets, respectively. Neural processes (NPs; (garnelo2018conditional; garnelo2018neural)) constitute a class of models that employ neural networks to learn a mapping $\eta:\bigcup_{k=0}^{\infty}(\mathcal{X}\times\mathcal{Y})^{k}\to\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$, which takes a finite (possibly empty) context set $\mathcal{D}_{c}$ and gives a stochastic process intended to approximate the Bayesian posterior induced by $\mu$. This approximation is constructed implicitly through its finite-dimensional marginals(garnelo2018neural; bruinsma2021gaussian; bruinsma2024convolutional; mathieu2023geometric). Denote ${\bm{x}}_{q}=(x_{q,k})_{k\in\mathcal{I}_{q}}$ and ${\bm{y}}_{q}=(y_{q,k})_{k\in\mathcal{I}_{q}}$. The NP approximation can then be expressed as
$q\!\left({\bm{y}}_{q}\mid{\bm{x}}_{q},\mathcal{D}_{c}\right)\approx p\!\left({\bm{y}}_{q}\mid{\bm{x}}_{q},,\mathcal{D}_{c}\right),$
where $q({\bm{y}}_{q}\mid{\bm{x}}_{q},\mathcal{D}_{c})$ and $p({\bm{y}}_{q}\mid{\bm{x}}_{q},\mathcal{D}_{c})$ are the finite-dimensional distributions—evaluated at ${\bm{x}}_{q}$—of the learned process $\eta[\mathcal{D}_{c}]$ and of the true posterior process induced by $\mu$, respectively.
In this work, we focus on conditional NPs (CNPs, garnelo2018conditional), which restrict $q$ to the family of mean-field Gaussians, i.e. $q\big({\bm{y}}_{q}\mid{\bm{x}}_{q},\mathcal{D}_{c}\big)=\prod_{k\in\mathcal{I}_{q}}q\!\left(y_{q,k}\mid x_{q,k},\mathcal{D}_{c}\right)$. In general, CNPs parameterization of $q\!\left(y_{q,k}\mid x_{q,k},\mathcal{D}_{c}\right)$ can be abstracted as a two-stage encoder-decoder pipeline (bruinsma2024convolutional; ashman2024translation; ashman2025gridded). The encoder $\varphi_{e}:\mathcal{X}\times\bigcup_{k=0}^{\infty}(\mathcal{X}\times\mathcal{Y})^{k}\rightarrow\mathcal{H}$ maps $x_{q,k}$ and $\mathcal{D}_{c}$ into a latent representation, while the decoder $\varphi_{d}:\mathcal{X}\times\mathcal{H}\rightarrow\Theta$ transforms $x_{q,k}$ and latent code into parameters $\theta\in\Theta$ of the predictive distribution $q_{\theta}(y_{q,k}\mid x_{q,k},\mathcal{D}_{c})$.
The vanilla CNP encoder summarizes the context set $\mathcal{D}_{c}$ using a permutation-invariant architecture (qi2017pointnet; zaheer2017deep). Each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$ is independently mapped by a shared feedforward network to a finite-dimensional embedding $\varepsilon_{c,k}$, which are subsequently aggregated—typically via mean pooling—into a single representation $\varepsilon_{c}$. Notably, this encoding is independent of query locations (i.e., $\varphi_{e}$ does not condition on ${\bm{x}}_{q}$). The decoder then fuses $\varepsilon_{c}$ with $x_{q,k}$ through another feedforward network to parameterize the Gaussian predictive distribution.
Although sum-pooling aggregation provides universal approximation guarantees (zaheer2017deep; bloem2020probabilistic), NPs employing such mechanisms often exhibit underfitting in practice (kim2019attentive). Prior works have partly attributed this phenomena to two primary factors(xu2020metafun): (1) the limitation of summaries with prespecified finite dimensionality in representing context sets of arbitrary size (wagstaff2019limitations), and (2) the shortcomings of simple sum or mean pooling operations to effectively capture rich interactions between context and query points(xu2020metafun; nguyen2022transformer).
Given that the problems tackled by NPs are inherently functional in nature, it is natural to seek embeddings that also exhibit a functional character. In this spirit, gordon2019convolutional formalize a general framework for constructing translation-equivariant prediction maps over sets, satisfying
$\eta\Big[\big\{(x_{c}+\tau,y_{c})\;\big|\;(x_{c},y_{c})\in\mathcal{D}_{c}\big\}\Big]\big((x_{q,k})_{k\in\mathcal{I}_{q}}\big)\;=\;\eta\big[\mathcal{D}_{c}\big]\big((x_{q,k}-\tau)_{k\in\mathcal{I}_{q}}\big),\vskip-2.84544pt$
for every $\tau\in\mathcal{X}$. They show that a broad family of such maps is expressed as $\eta[\mathcal{D}_{c}]=\varphi_{d}\big[\varphi_{e}\left[\mathcal{D}_{c}\right]\big]$, where the functional embedding is defined by
$\varphi_{e}[\mathcal{D}_{c}](\cdot)=\sum_{(x_{c},\,y_{c})\in\mathcal{D}_{c}}\phi(y_{c})\,\psi_{e}(\cdot-x_{c}),\vskip-2.84544pt$ (3)
and $\varphi_{d}:\mathcal{H}\to C_{b}(\mathcal{X},\mathcal{Y})$ is a translation-equivariant decoder operating on a function space $\mathcal{H}$. Here, $C_{b}(\mathcal{X},\mathcal{Y})$ denotes the space of bounded continuous functions from $\mathcal{X}$ to $\mathcal{Y}$; $\phi(y)=(1,y)$ [^3] For most applications, this representation suffices. Generally, $\phi(y)=(1,y,\dots,y^{M})$, where $M$ is the multiplicity of repeated inputs. See gordon2019convolutional and bruinsma2024convolutional for further discussion. and $\psi_{e}:\mathcal{X}\to{\mathbb{R}}$ is a continuous, strictly positive-definite kernel typically chosen to be a Gaussian.
In implementing ConvCNPs, $\varphi_{e}[\mathcal{D}_{c}]$ is first evaluated on a uniform grid $\mathcal{G}\subset\mathcal{X}$ that spans the joint support of both ${\bm{x}}_{q}$ and ${\bm{x}}_{c}$. This yields a discretized representation $(\varphi_{e}[\mathcal{D}_{c}](x_{g}))_{x_{g}\in\mathcal{G}}$, which is then passed through $\varphi_{d}$ to produce $\big(\varphi_{d}\big[\varphi_{e}[\mathcal{D}_{c}]\big](x_{g})\big)_{x_{g}\in\mathcal{G}}$. Because ${\bm{x}}_{q}$ may not lie exactly on $\mathcal{G}$, ConvCNPs employ an additional interpolation step to obtain smooth predictions at ${\bm{x}}_{q}$. Specifically
$(\theta_{q,k})_{k\in\mathcal{I}_{q}}=\bigg(\,\sum_{x_{g}\in\mathcal{G}}\varphi_{d}\big[\varphi_{e}[\mathcal{D}_{c}]\big](x_{g})\,\psi_{d}(x_{q,k}-x_{g})\bigg)_{k\in\mathcal{I}_{q}},\vskip-2.84544pt$ (4)
@@ 2.2 Neural Processes
where $\psi_{d}$ is another strictly positive kernel. Note that this step can be viewed as part of the decoder itself, thereby preserving the overall encoder–decoder abstraction discussed earlier.

### v6

@@ 1 Introduction
Stochastic processes offer a mathematical framework for modeling systems that evolve with inherent randomness over continuous domains such as time and space. They underpin a wide range of scientific applications—from spatio-temporal climate dynamics to biological and physical systems—thereby motivating the development of machine learning methods that can learn from data generated by such phenomena (Mathieu et al., 2021; Vaughan et al., 2021; Allen et al., 2025; Ashman et al., 2025; Dupont et al., 2021). Among classical approaches, Gaussian processes (GPs; Rasmussen et al., 2006) provide a Bayesian framework with closed-form inference and uncertainty quantification. However, their cubic computational cost from matrix inversion and the difficulty of specifying suitable kernels—especially in high-dimensional settings—limit their scalability.
@@
Motivated by the success of deep neural networks in large-scale function approximation, neural network-based alternatives have emerged. Neural processes (NPs; Garnelo et al., 2018a, b) exemplify this paradigm, combining ideas from GPs and deep learning within a meta-learning framework. By exposing the model to multiple realizations of an underlying stochastic process, each treated as a distinct task, NPs learn shared structures across tasks to parameterize a neural mapping that directly gives an approximation of the corresponding posterior predictive distribution (Bruinsma, 2024). Once trained, the model enables efficient probabilistic predictions on new tasks without further training.
@@
Since the introduction of conditional neural processes (CNPs, Garnelo et al. (2018a)) as the first class within the NPs family (Jha et al., 2022), numerous extensions have been proposed to enhance their effectiveness. One prominent line of work focuses on incorporating explicit inductive biases into CNPs in order to better capture the symmetries that commonly arise in scientific applications (Gordon et al., 2019; Kawano et al., 2021; Holderrieth et al., 2021; Huang et al., 2023; Ashman et al., 2024a, b). Another major direction seeks to move beyond the mean-field factorized Gaussian predictive distributions to which CNPs are limited. A widely adopted approach augments CNPs with stochastic latent variables, giving rise to the family of latent neural processes (LNPs; Garnelo et al. (2018b); Louizos et al. (2019); Wang and Van Hoof (2020); Foong et al. (2020); Lee et al. (2020); Volpp et al. (2021); Wang et al. (2022); Wang and van Hoof (2022); Kim et al. (2022); Jung and Park (2023); Lee et al. (2023); Xu et al. (2023)). Complementary efforts explore autoregressive prediction schemes (Bruinsma et al., 2023; Nguyen and Grover, 2022), Gaussian predictive distributions with non-diagonal covariances (Bruinsma et al., 2021; Markou et al., 2022), and quantile-based parameterizations of the predictive distribution (Mohseni et al., 2023).
@@
This work focuses on CNPs, particularly convolutional CNPs (ConvCNPs; Gordon et al. (2019)), which were the first to endow NPs with translation equivariance. ConvCNPs introduce the convolutional deep set construction which characterizes a broad class of translation-equivariant mappings over finite, potentially unstructured sets of observations as a composition of a functional embedding with a translation-equivariant operator, typically realized through convolutional neural networks (CNNs; Fukushima (1980); LeCun et al. (1989, 1998)).
@@
Despite their effectiveness, ConvCNPs can struggle to aggregate information from observations spread across large spatial domains—a challenge that becomes particularly pronounced in sparse data regimes. This limitation stems from their reliance on local convolutional kernels with small receptive fields, which hampers their ability to model long-range dependencies. A natural remedy is to enlarge the kernel size to extend the receptive field; however, this approach rapidly increases the number of model parameters and computational cost (Romero et al., 2021). Alternatively, transformer-based architectures can capture long-range interactions but incur quadratic, rather than linear, computational complexity in the number of observations (Vaswani et al., 2017; Nguyen and Grover, 2022).
@@
In this work, we pursue an alternative paradigm that represents functions in the frequency domain, inspired by the well-established observation that many natural processes exhibit energy concentration in low-frequency bands (Field, 1987; Ruderman and Bialek, 1993; Wainwright and Simoncelli, 1999). This property allows for efficient approximation using only a subset of dominant spectral coefficients, enabling tractable computations while preserving the signal’s global structure. By parameterizing convolution kernels directly in the Fourier domain over a finite set of frequencies and leveraging the convolution theorem, we can attain large effective receptive fields without incurring prohibitive computational costs.
@@
While spectral methods have been extensively studied in neural operator learning for partial differential equations (PDEs, Chen and Chen (1995); Kovachki et al. (2021, 2023)), their application within NPs framework remains relatively unexplored. To bridge this gap, we propose spectral convolutional conditional neural processes (SConvCNPs)—a class of models that adopt Fourier neural operators (FNOs; Li et al. (2020a)) to realize global spectral convolution while maintaining computational efficiency. Across a suite of synthetic and real-world benchmarks, SConvCNPs perform competitively with state-of-the-art baselines, illustrating how ideas from neural operators can enhance the flexibility and performance of NPs.
@@ 2.1 Fourier Neural Operators
Neural operators (Chen and Chen, 1995; Li et al., 2020a; Kovachki et al., 2023) are neural network architectures designed to learn mappings between function spaces rather than finite-dimensional vectors. Like conventional feed-forward networks, they consist of stacked layers that alternate between operator-based transformations and pointwise nonlinearities. A common transformation is an integral operator with kernel $\kappa:\mathcal{X}\times\mathcal{X}\to\mathcal{Y}$, acting on an input function $v:\mathcal{X}\to\mathcal{Y}$ as
$\mathcal{K}[v](x)=\int\kappa(x,s)\,v(s)\,\mathrm{d}s,$
@@
where $\mathcal{X}={\mathbb{R}}^{d_{x}}$ and $\mathcal{Y}={\mathbb{R}}^{d_{y}}$ are Euclidean spaces with $d_{x},d_{y}\in\mathbb{N}$. While this work focuses on operators linear in $v$, nonlinear formulations have also been explored, including continuous formulations of softmax attention (Ashman et al., 2024a; Calvello et al., 2024). When the kernel is stationary—that is, $\kappa(x,s)=\kappa(x-s)$—the operator reduces to a convolution, $\mathcal{K}[v]=\kappa\ast v$, connecting neural operators to CNNs where $\kappa$ is parameterized by learnable weights.
@@
Convolutional kernels are typically spatially local with limited receptive fields (Luo et al., 2016; Peng et al., 2017; Wang et al., 2018). Modeling long-range dependencies thus requires large kernels, substantially increasing parameter count. The Fourier Neural Operator (FNO; Li et al. (2020a)) addresses this limitation by exploiting the convolution theorem (Bracewell and Kahn, 1966; Oppenheim, 1999), which expresses convolution as
$\mathcal{K}[v](x)=\mathcal{F}^{-1}\Big[\widehat{\kappa}(\xi)\cdot\widehat{v}(\xi)\Big](x),$ (1)
@@
where $\widehat{f}:=\mathcal{F}[f]$ denotes the Fourier transform of $f$, and $\mathcal{F}^{-1}$ denotes the inverse Fourier transform. Rather than parameterizing $\kappa$ in the spatial domain, the FNO learns its Fourier representation $\widehat{\kappa}$ directly. When $v$ is approximately band-limited—i.e., $\widehat{v}(\xi)$ has negligible energy for $\|\xi\|>\xi_{0}$—high-frequency components can be truncated with minimal information loss. This property, observed in many natural signals (Field, 1987; Ruderman and Bialek, 1993; Wainwright and Simoncelli, 1999), allows setting $\widehat{\kappa}(\xi)=0$ outside the retained band without significant degradation.
In practice, functions are accessible only through discrete samples, requiring the discrete Fourier transform (DFT) for domain transitions. Given samples of $v$ on a uniform grid $\mathcal{G}\subset\mathcal{X}$, the FNO applies the DFT via the fast Fourier transform (FFT; Cooley and Tukey (1965); Frigo and Johnson (2005)). The resulting spectrum is truncated to a finite set of frequency modes $\Xi\subset\mathbb{R}^{d_{x}}$ assumed to capture most of the signal’s energy. For each retained frequency $\xi\in\Xi$, the Fourier kernel is parameterized by learnable complex weights $W_{\xi}\in\mathbb{C}^{d_{y}}$, so that $\widehat{\kappa}(\xi)=W_{\xi}$. [^1] More generally, a matrix-valued parameterization $W_{\xi}\in\mathbb{C}^{c_{\text{out}}\times d_{y}}$ is used, where $c_{\text{out}}$ is the number of output channels, enabling joint mixing of input channels and projection to different dimensionality. This formulation implicitly imposes spatial periodicity, as the kernel is represented using discrete harmonics (equivalently, a Dirac comb in frequency space). After pointwise multiplication in frequency space, an inverse FFT returns the operator output to the spatial domain.
2.2 Neural Processes
Let $\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$ denote the space of probability measures over measurable functions $f:\mathcal{X}\to\mathcal{Y}$, which we interpret as stochastic processes indexed by $\mathcal{X}$. We assume an unknown data-generating process $\mu\in\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$, from which latent functions are drawn. A task consists of a finite collection of noisy input–output observations generated from a realization $f\sim\mu$, partitioned into a context set and a query set:
$\mathcal{D}=(\mathcal{D}_{c},\mathcal{D}_{q}),\qquad\mathcal{D}_{c}=\{(x_{c,k},y_{c,k})\}_{k=1}^{n_{c}},\qquad\mathcal{D}_{q}=\{(x_{q,l},y_{q,l})\}_{l=1}^{n_{q}}.$ (2)
Observations are generated according to
$y_{c,k}=f(x_{c,k})+\epsilon_{c,k},\qquad y_{q,l}=f(x_{q,l})+\epsilon_{q,l},$
where $\epsilon_{c,k},\epsilon_{q,l}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma_{0}^{2})$ and $\sigma_{0}>0$ denotes the observation noise scale.
Neural processes (NPs; (Garnelo et al., 2018a, b)) constitute a class of models that use neural networks to learn a mapping $\eta:\mathcal{S}(\mathcal{X}\times\mathcal{Y})\to\mathcal{P}(\mathcal{Y}^{\mathcal{X}})$, where $\mathcal{S}(\mathcal{X}\times\mathcal{Y})$ denotes the collection of all finite subsets of $\mathcal{X}\times\mathcal{Y}$. Given a context set $\mathcal{D}_{c}$, $\eta$ outputs a stochastic process intended to approximate the Bayesian posterior over functions induced by $\mu$ and conditioned on $\mathcal{D}_{c}$. Typically, this process is specified implicitly through its finite-dimensional marginals (Garnelo et al., 2018b; Bruinsma et al., 2021; Bruinsma, 2024; Mathieu et al., 2023). Let $\mathbf{x}_{q}=(x_{q,l})_{l=1}^{n_{q}}$ and $\mathbf{y}_{q}=(y_{q,l})_{l=1}^{n_{q}}$. We denote by $\eta[\mathcal{D}_{c};\mathbf{x}_{q}](\cdot)$ the finite-dimensional distribution of the process $\eta[\mathcal{D}_{c}]$ evaluated at $\mathbf{x}_{q}$, and by $\mu_{\mathbf{x}_{q}}(\cdot\mid\mathcal{D}_{c})$ the corresponding finite-dimensional marginal of the true posterior process. Informally, the NP approximation aims to satisfy
$\eta[\mathcal{D}_{c};\mathbf{x}_{q}](\cdot)\;\approx\;\mu_{\mathbf{x}_{q}}(\cdot\mid\mathcal{D}_{c}),$
for all $\mathbf{x}_{q}$ and $\mathcal{D}_{c}$ (Bruinsma et al., 2021; Bruinsma, 2024). With slight abuse of notation, we write $\eta[\mathcal{D}_{c};\mathbf{x}_{q}](\mathbf{y}_{q})$ for the density of this finite-dimensional distribution evaluated at $\mathbf{y}_{q}$.
In this work, we focus on conditional neural processes (CNPs; Garnelo et al. (2018a)), which restrict these finite-dimensional distributions to mean-field Gaussians, i.e. $\eta[\mathcal{D}_{c};\mathbf{x}_{q}]=\prod_{l=1}^{n_{q}}\eta[\mathcal{D}_{c};x_{q,l}]$, where each marginal $\eta[\mathcal{D}_{c};x_{q,l}]$ is Gaussian. Each predictive marginal is typically parameterized via a two-stage encoder–decoder architecture (Bruinsma, 2024; Ashman et al., 2024a, 2025). The encoder $\varphi_{e}:\mathcal{S}(\mathcal{X}\times\mathcal{Y})\to\mathcal{H}$ maps the context set $\mathcal{D}_{c}$ to a latent representation, while the decoder $\varphi_{d}:\mathcal{H}\to\Theta^{\mathcal{X}}$ maps this representation to a function that assigns, to each query $x_{q,l}$, parameters $\theta(x_{q,l})\in\Theta$ of the predictive distribution $\eta[\mathcal{D}_{c};x_{q,l}]$.
The vanilla CNP encoder summarizes $\mathcal{D}_{c}$ using a permutation-invariant architecture (Qi et al., 2017; Zaheer et al., 2017). Each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$ is independently mapped by a shared network to a finite-dimensional embedding $\varepsilon_{c,k}$, which are then aggregated—typically via mean pooling—into a single representation $\varepsilon_{c}$. Notably, this encoding is independent of the query locations. The decoder then combines $\varepsilon_{c}$ with each $x_{q,l}$ to parameterize the Gaussian predictive distribution.
Although sum-pooling aggregation provides universal approximation guarantees (Zaheer et al., 2017; Bloem-Reddy and Teh, 2020), NPs employing such mechanisms often exhibit underfitting in practice (Kim et al., 2019). Prior works have partly attributed this phenomena to two primary factors (Xu et al., 2020): (1) the limitation of summaries with prespecified finite dimensionality in representing context sets of arbitrary size (Wagstaff et al., 2019), and (2) the shortcomings of simple sum or mean pooling operations to effectively capture rich dependencies between different points (Xu et al., 2020; Nguyen and Grover, 2022).
Since NPs address inherently functional learning problems, it is natural to consider embeddings that themselves take the form of functions. Gordon et al. (2019) introduce a framework for translation-equivariant prediction maps over sets, satisfying
$\eta\Big[\{(x_{c}+\tau,y_{c})\mid(x_{c},y_{c})\in\mathcal{D}_{c}\};\mathbf{x}_{q}\Big]=\eta\big[\mathcal{D}_{c};\mathbf{x}_{q}-\tau\big],$
for all translations $\tau$, context sets $\mathcal{D}_{c}$, and collections of query locations $\mathbf{x}_{q}$, where subtraction by $\tau$ acts pointwise, i.e., $\mathbf{x}_{q}-\tau=(x_{q,l}-\tau)_{l=1}^{n_{q}}$. They show that a broad class of such maps can be written as $\eta[\mathcal{D}_{c}]=\varphi_{d}\circ\varphi_{e}[\mathcal{D}_{c}]$, where the functional embedding is defined by
$\varphi_{e}[\mathcal{D}_{c}](x)=\sum_{(x_{c},y_{c})\in\mathcal{D}_{c}}\!\!\!\!\phi(y_{c})\,\psi_{e}(x-x_{c}).$ (3)
@@ 2.2 Neural Processes
In convolutional conditional neural processes (ConvCNPs), the functional embedding $\varphi_{e}[\mathcal{D}_{c}]$ is evaluated on a uniform grid $\mathcal{G}\subset\mathcal{X}$ covering the joint support of $\mathbf{x}_{c}$ and $\mathbf{x}_{q}$, yielding the discretized representation $(\varphi_{e}[\mathcal{D}_{c}](x_{g}))_{x_{g}\in\mathcal{G}}$. This representation is processed by $\varphi_{d}$, and predictions at query locations are obtained via kernel interpolation:
$\theta_{q,l}=\sum_{x_{g}\in\mathcal{G}}(\varphi_{d}\circ\varphi_{e})[\mathcal{D}_{c}](x_{g})\,\psi_{d}(x_{q,l}-x_{g}),\qquad l=1,\dots,n_{q},$ (4)
@@
where $\psi_{d}$ is another strictly positive-definite kernel. This interpolation step may be viewed as part of the decoder, preserving the overall encoder–decoder abstraction.